OpenStudy (anonymous):

HELP PLEASE! Which equation can be used to solve the following word problem? Emilio has 6 more nickels than dimes, and the total value of his coins is $2.10. How many nickels does he have? Let n represent the number of nickels.

OpenStudy (anonymous):

A. 0.10n + 0.05n + 6 = 2.10 B. 0.10n + 0.05(n + 6) = 2.10 C. 0.10(n + 6) + 0.05n = 2.10 D. 0.10(n – 6) + 0.05n = 2.10

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

I thought it would be: $0.10+.30+$0.05=$2.10 but i was wrong

jimthompson5910 (jim_thompson5910):

Hint dimes = nickels + 6 d = n + 6 n = d - 6

OpenStudy (anonymous):

so... .10+.05(n+6)?

jimthompson5910 (jim_thompson5910):

no you don't have n dimes, you have n-6 dimes

OpenStudy (anonymous):

answer is C

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

thank you. you didn't have to tell me tho. but thanks anyways

jimthompson5910 (jim_thompson5910):

it's not C

jimthompson5910 (jim_thompson5910):

you don't have n+6 dimes, you have n-6 dimes

OpenStudy (anonymous):

oh so D

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

oh wait, had them switched, my bad nickels = dimes + 6 n = d+6 d = n - 6 but the answer is still D

OpenStudy (anonymous):

ok thanks :)

jimthompson5910 (jim_thompson5910):

np